OpenStudy (anonymous):

Is x/6y * 3y^2/11 equal to xy/22, y ≠ 0

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

they are equal I said...

OpenStudy (anonymous):

Ok =)

OpenStudy (anonymous):

yes x/6y * 3y^2/11 = x*3y^2/6y*11 = 3xy^2/66y = xy/22

OpenStudy (anonymous):